Two bodies have a mass of ratio 3 is to 4 the force applied on bigger mass production and acceleration of 12 mass per second what would be the acceleration of other body to the same force acts on it
1
Expert's answer
2020-07-10T10:26:47-0400

The ratio of masses is "\dfrac{m_1}{m_2} = \dfrac{3}{4}". According to the second Newton's law, the force is:


"F = ma"

where "a" is acceleration delivered to the body.

If the force is the same for both bodies, then:


"m_1a_1 = m_2a_2\\
\dfrac{m_1}{m_2} = \dfrac{a_2}{a_1} = \dfrac{3}{4}"

From the last equation:


"a_2 = \dfrac{3}{4}a_1"

If "a_1 = 12 m\/s^2", then:


"a_2 = \dfrac{3}{4}12m/s^2 = 9m/s^2"

Answer. 9 m/s^2.